# Learning Partial Graph Matching via Optimal Partial Transport

### v5

@@ 7 Partial Graph matching Architecture
The matching biases $\alpha$ and $\beta$ are calculated based on the cross-graph node-to-node affinity matrix $A$. For each node $i\in V_{S}$, define $r_{i}=\max_{1\leq j\leq n}A_{ij}$, which represents the highest affinity of node $i$ in $\mathcal{G}_{S}$ with any node in $\mathcal{G}_{T}$. Similarly, for each node $j\in V_{T}$, define $r_{j}=\max_{1\leq i\leq m}A_{ij}$, representing the highest affinity of node $j$ in $\mathcal{G}_{T}$ with any node in $\mathcal{G}_{S}$. The matching bias is then calculated as $\alpha_{i}=\sigma(w_{rs}\times r_{i})$ for each $i\in V_{S}$ and $\beta_{j}=\sigma(w_{rs}\times r_{j})$ for each $j\in V_{T}$, where $\sigma(\cdot)$ is the sigmoid function, and $w_{rs}$ is a learnable parameter. A value of $\alpha_{i}$ closer to 1 indicates a higher chance of node $i$ being matched, while a value closer to 0 indicates a lower chance. Similarly, for $\beta_{j}$, a value closer to 1 suggests a higher chance of node $j$ being matched, and a value closer to 0 suggests a lower chance.
@@ 8 Experiments
In this section, we evaluate our proposed partial graph matching approach to empirically verify its performance and robustness, addressing the following aspects: (1) We first evaluate the efficacy of our approach through experiments on image keypoint matching datasets and Protein-Protein Interaction (PPI) networks under varying noise levels; (2) We analyze the effects of the matching biases $\alpha$ and $\beta$ on overall performance by considering two model variants: OPGM, with fixed equal matching biases ($\alpha=\mathbf{1}_{m},\beta=\mathbf{1}_{n}$), and OPGM-rs, with learnable matching biases. Note that when training OPGM, $L=L_{cost}$, as we do not need to learn matching biases. (3) We further conduct a sensitivity analysis on the unbalancedness parameter $\rho$ and regularization parameter $\lambda$ to understand their impact on partial graph matching performance; (4) We analyze the runtime efficiency of our approach compared to other baselines. Our code is available at GitHub: https://github.com/Gathika94/OPGM.git.
@@ 8.1 Experimental Setup
GCAN OPGM 51.9 43.8 66.6 28.9 60.9 60.6 37.8 67.8 37.7 64.3 58.9 47.6 47.8 43.1 77.3 49.5 82.1 90.9 56.5 ± 0.4
GCAN OPGM-rs 52.5 43.2 67.9 32.3 61.2 61.2 39.9 68.9 38.2 67.1 60.0 50.1 47.4 46.3 78.8 50.2 82.9 92.7 57.8 ± 0.2
@@
Dataset Name IMCPT 50 IMCPT 100 GM Network PMH reichstag sacre st peters mean reichstag sacre st peters mean NGM-v2 dummy 88.5 56.1 63.0 69.2±0.5 80.0 57.0 71.3 69.5±0.3 NGM-v2 AFAT-U 90.5 58.7 66.9 72.0±0.3 81.7 57.0 72.2 70.3±0.2 NGM-v2 AFAT-I 92.3 58.7 66.7 72.8±0.4 82.0 57.0 71.4 70.1±0.3 GCAN ILP 87.2 55.1 63.0 68.4±0.5 80.4 55.7 72.8 69.6±0.4 GCAN AFAT-U 86.9 59.4 67.1 71.1±0.4 82.6 58.2 73.8 71.5±0.2 GCAN AFAT-I 91.0 60.3 67.3 72.9±0.6 82.7 57.8 72.4 70.9±0.4 GCAN OPGM 91.2 57.0 68.1 71.9±0.2 81.7 59.1 76.1 72.3±0.4 GCAN OPGM-rs 91.7 59.2 67.2 72.7±0.1 82.6 59.6 75.9 72.7±0.2

### v6

@@ 7 Partial Graph matching Architecture
The matching biases $\alpha$ and $\beta$ are calculated based on the cross-graph node-to-node affinity matrix $A$. Let $A^{+}\in\mathbb{R}_{\geq 0}^{m\times n}$ be defined as $A^{+}_{ij}=\max\{A_{ij},0\}$. For each node $i\in V_{S}$, define $r_{i}=\max_{1\leq j\leq n}A^{+}_{ij}$. Similarly, for each node $j\in V_{T}$, define $r_{j}=\max_{1\leq i\leq m}A^{+}_{ij}$. The matching bias is then calculated as $\alpha_{i}=2\times(\sigma(w_{rs}\times r_{i})-0.5)$ for each $i\in V_{S}$ and $\beta_{j}=2\times(\sigma(w_{rs}\times r_{j})-0.5)$ for each $j\in V_{T}$, where $\sigma(\cdot)$ is the sigmoid function, and $w_{rs}\geq 0$ is a learnable parameter. A value of $\alpha_{i}$ or $\beta_{j}$ closer to 1 indicates a higher chance of node $i$ or $j$ being matched, while a value closer to 0 indicates a lower chance. Similarly, for $\beta_{j}$, a value closer to 1 suggests a higher chance of node $j$ being matched, and a value closer to 0 suggests a lower chance.
@@ 8 Experiments
In this section, we evaluate our proposed partial graph matching approach to empirically verify its performance and robustness, addressing the following aspects: (1) We first evaluate the efficacy of our approach through experiments on image keypoint matching datasets and Protein-Protein Interaction (PPI) networks under varying noise levels; (2) We analyze the effects of the matching biases $\alpha$ and $\beta$ on overall performance by considering two model variants: OPGM, with fixed equal matching biases ($\alpha=\mathbf{1}_{m},\beta=\mathbf{1}_{n}$), and OPGM-rs, with learnable matching biases. Note that when training OPGM, $L=L_{cost}$, as we do not need to learn matching biases. (3) We further conduct a sensitivity analysis on the unbalancedness parameter $\rho$ and regularization parameter $\lambda$ to understand their impact on partial graph matching performance; (4) We analyze the runtime efficiency of our approach compared to other baselines. Our code is available at GitHub: https://github.com/Gathika94/OPGMrs.git.
@@ 8.1 Experimental Setup
GCAN OPGM 51.9 43.8 66.6 28.9 60.9 60.6 37.8 67.8 37.7 64.3 58.9 47.6 47.8 43.1 77.3 49.5 82.1 90.9 56.5 $\pm$ 0.4
GCAN OPGM-rs 53.0 43.5 66.7 32.1 61.7 61.4 40.7 68.8 38.5 65.8 59.5 51.1 47.2 46.2 78.6 48.7 83.3 92.3 57.7 $\pm$ 0.2
@@
Dataset Name IMCPT 50 IMCPT 100 GM Network PMH reichstag sacre st peters mean reichstag sacre st peters mean NGM-v2 dummy 88.5 56.1 63.0 69.2±0.5 80.0 57.0 71.3 69.5±0.3 NGM-v2 AFAT-U 90.5 58.7 66.9 72.0±0.3 81.7 57.0 72.2 70.3±0.2 NGM-v2 AFAT-I 92.3 58.7 66.7 72.8±0.4 82.0 57.0 71.4 70.1±0.3 GCAN ILP 87.2 55.1 63.0 68.4±0.5 80.4 55.7 72.8 69.6±0.4 GCAN AFAT-U 86.9 59.4 67.1 71.1±0.4 82.6 58.2 73.8 71.5±0.2 GCAN AFAT-I 91.0 60.3 67.3 72.9±0.6 82.7 57.8 72.4 70.9±0.4 GCAN OPGM 91.2 57.0 68.1 72.1±0.2 81.7 59.1 76.1 72.3±0.4 GCAN OPGM-rs 91.9 59.2 67.0 72.7±0.1 82.3 60.5 75.7 72.8±0.1